OpenStudy (anonymous):

i got a few questions Q1) the origin and has a gradient of 2 Q2) -3y+9x

OpenStudy (lochana):

Q1) y = 2x

OpenStudy (anonymous):

ahh this is Coordinate geometry i dont get what does origin means

OpenStudy (agreene):

origin is where the axes cross, so the point (0,0)

OpenStudy (anonymous):

ohhh okey so the formula will be y-0=m(x-0)

OpenStudy (agreene):

yeah, except they tell you m (slope) the word gradient means the same thing.

OpenStudy (anonymous):

but it say m=2 so it will be y-0=2(x-0) is the answer 0=2x-y-0

OpenStudy (agreene):

y=2x would be the normal form, but yea it could also be 0=2x-y ... i'd put y=2x though.

OpenStudy (anonymous):

thanks

OpenStudy (agreene):

The second question, I'm not sure what you're asking exactly.

OpenStudy (anonymous):

Well the second question is to find the graidient-intercept form and i dont get it

OpenStudy (agreene):

like you mentioned, the normal form is y=mx+b where m is slope and b is the x-intercept. So, what it's asking is for you to put -3y+9x=0 in that form. (solve for y)

OpenStudy (anonymous):

isnt x have to be first then y

OpenStudy (anonymous):

cause my friend helped me with most of it except that question and he wrote on my work 0=ax+by+c

OpenStudy (agreene):

that form isn't "gradient-intercept" form. That form is defined as being y=mx+b As I mentioned we need to solve for y; we would start by isolating that variable... -3y+9x=0 subtract 9x from both sides and then we have -3y=-9x